OpenStudy (anonymous):

lim x approaching 1 (x^2-1)/(x^2-x)

OpenStudy (freckles):

factor factor!

OpenStudy (freckles):

x^2-1 is a difference of squares and x^2-x you have each term having a common factor of x

OpenStudy (anonymous):

still dont know what to do

OpenStudy (freckles):

do you know how to factor x^2-1?

OpenStudy (freckles):

x^2-a^2=(x-a)(x+a) and you have a=1 so x^2-1=?

OpenStudy (anonymous):

0

OpenStudy (freckles):

x^2-1 is not always 0

OpenStudy (freckles):

x^2-1=(x-1)(x+1) just using the formula I gave you above

OpenStudy (freckles):

x^2-x=x(x-1)

OpenStudy (freckles):

\[\lim_{x \rightarrow 1} \frac{x^2-1}{x^2-x}=\lim_{x \rightarrow 1} \frac{(x-1)(x+1)}{x(x-1)}\] you see anything that cancels?

OpenStudy (anonymous):

wow okay i wrote it wrong , thanks